Please help me Perform the indicated operation. Thanks
4x5-3x4^2/6-8x2

Answer by ptaylor(1894)

The problem:
4x5-3x4^2/6-8x2
OK, I think that basically we are required to simplify the expression but I'm not exactly sure what the expression is:
is it:
(4)x(5)-(((3)x(4^2))/6)-(8)x(2) if so, lets take each term by itself and simplify it, then we'll put the terms back together:
(4)x(5)=20
4^2=16
(3)x(4^2)=48
48/6=8
(8)x(2)=16
Putting it back together, we have:
20-8-16=20-24=-4

or is the problem this:
((4)x(5)-(((3)x(4^2))/(6-(8)x(2)) If so, lets get the individual terms that we have already generated and put them back using this layout:
(20-48)/(6-16)=
-28/-10=14/5
Hope this helps-----ptaylor